OpenStudy (anonymous):

Find the fundamental period of f when f(x) = 2 cos(9x) .

OpenStudy (jamesj):

What is the period of cos x?

OpenStudy (anonymous):

pi

OpenStudy (jamesj):

Given that, if we write T for the period of your function, for what value of T is f(t+T) = f(t) i.e., 2cos(9(t+T)) = 2cos(9t) i.e., cos(9t + 9T) = cos(9t) It must be that 9T = 2pi Hence the period of your function f, is T = 2pi/9.

OpenStudy (jamesj):

The period of cos x and sin x is 2pi. That is why we set 9T = 2pi.